OpenStudy (anonymous):

Solve: 4(x + 6) = 6(x - 2)

OpenStudy (anonymous):

how do you expand 4(x+6) ?

OpenStudy (anonymous):

4x + 24 = 6x - 12 24 + 12 = 6x - 4x 36 = 2x 18 = x

OpenStudy (anonymous):

let askers try it by themselves

OpenStudy (anonymous):

I just showed her how to do it step by step so that she could do the rest by herself.

OpenStudy (anonymous):

You know I think i have it thanks for explaining it @JBrightman3 , I just needed to see it in front of me..an it's all coming back to me :)

OpenStudy (anonymous):

No problem studmuffin. Good luck!

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

@JBrightman3 is correct